OpenStudy (diyadiya):

If \( a(p+q)^2+2bpq+c=0\) and \(a(p+r)^2+2bpr+c=0\) then qr = ?

OpenStudy (anonymous):

can't wait to see this!

OpenStudy (diyadiya):

lol

OpenStudy (anonymous):

This is a perfect candidate for my problem of the day.

OpenStudy (anonymous):

Good one diya!

OpenStudy (diyadiya):

lol So how do i do it ?

OpenStudy (anonymous):

Lol do you think I can answer everything just at sight? :P

OpenStudy (diyadiya):

i did :)

OpenStudy (anonymous):

Is it \(\huge \frac{(ap^2 + c) }{ a} \) ?

OpenStudy (diyadiya):

Yup =D

OpenStudy (anonymous):

gimmick is...?

OpenStudy (anonymous):

Wo :D. It wasn't that hard as it look like :)

OpenStudy (anonymous):

\[ a(p+q)^2 + 2bpq + c = 0 \implies aq^2 + 2p(a+b)q + ap^2 + c = 0 \] and \[ a(p+r)^2 +2bpr + c = 0 \implies ar^2 + 2p(a+b)r + ap^2 + c = 0 \] can you see the rest? ;)

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

I agree sat :)

OpenStudy (diyadiya):

ok let me see..

OpenStudy (anonymous):

Don't see, observe :D I have to go now. It's nice solving your problems :)

OpenStudy (diyadiya):

\[aq^2+2p(a+b)q=ar^2+2p(a+b)r\]

OpenStudy (anonymous):

\(q\) and \(r\) are the equation of \[ ax^2 + 2p(a+b)x + ap^2 + c = 0 \]

OpenStudy (diyadiya):

okay!

OpenStudy (diyadiya):

okay got it Thanks!!!

OpenStudy (anonymous):

You are welcome, but I haven't done anything special.